Accounts receivable analysis

The following data are taken from the financial statements of Sigmon Inc. Terms of all sales are 2/10, n/45.

20Y3 20Y2 20Y1
Accounts receivable, end of year $720,000 $650,000 $585,000
Sales on account 5,627,000 4,531,000
  1. For 20Y2 and 20Y3, determine (1) the accounts receivable turnover and (2) the number of days' sales in receivables. Assume a 365-day year. Do not round intermediate calculations. Round your answers to one decimal place.

    20Y3 20Y2
    1. Accounts receivable turnover
    2. Number of days' sales in receivables days days
  2. The collection of accounts receivable has (deteriorated or improved) This can be seen in both the (Decrease or Increase) in accounts receivable turnover and the (Decrease or Increase) in the collection period.
Answer #1

Answer

  • Working

Working

20Y3

20Y2

A

Beginning Accounts receivables

$          650,000.00

$        585,000.00

B

Ending Accounts receivables

$          720,000.00

$        650,000.00

C = A+B

Total

$      1,370,000.00

$     1,235,000.00

D = C/2

Average Accounts receivables

$          685,000.00

$        617,500.00

E

Net sales on account

$      5,627,000.00

$     4,531,000.00

F = E/D

Accounts receivables Turnover

8.21459854

7.337651822

Working

20Y3

20Y2

A

No. of Days

365

365

B

Accounts receivables Turnover

8.21459854

7.337651822

C =A/B

No. of days sales in receivables

44.4

49.7

  • Answers

20Y3

20Y2

1. Accounts receivable turnover

8.2

times

7.3

times

2. Number of days' sales in receivables

44.4

Days or 44.5 days

49.7

Days or 50 days

  • Requirements

The collection of accounts receivable has improved. This can be seen in both the Increase in accounts receivable turnover and the Decrease in the collection period.